Bobet is 3 more than 4 times the age of his son. Next year The sum of their ages will be 40. How old is Bobet?


Found 2 solutions by ankor@dixie-net.com, jorel555:
Answer by ankor@dixie-net.com(22740)

Bobet is 3 more than 4 times the age of his son.
b = 4s + 3
Next year The sum of their ages will be 40.
b + 1 + s + 1 = 40
b + s + 2 = 40
b + s = 38
replace b with (4s+3) from the 1st statement
(4s+3) + s = 38
4s + s = 38 - 3
5s = 35
s = 35/5
s = 7 yrs is the son's age
:
How old is Bobet?
b = 4(7) + 3
b = 31 is Bob's age
:
:
Check solutions in the statement
"Next year The sum of their ages will be 40".
31 + 1 + 7 + 1 = 40

Answer by jorel555(1290)

Let n be the age of Bobet's son. Then Bobet is 4n+3. The next year, their ages will be n+1 and 4n+4. So:
n+1+4n+4=40
5n=35
n=7
Bobet's son is now 7; Bobet is 31. ☺☺☺☺